### v2

# Rapid Adoption, Hidden Risks: The Dual Impact of Large Language Model Customization

Abstract
The increasing demand for customized Large Language Models (LLMs) has led to the development of solutions like GPTs. These solutions facilitate tailored LLM creation via natural language prompts without coding. However, the trustworthiness of third-party custom versions of LLMs remains an essential concern. In this paper, we propose the first instruction backdoor attacks against applications integrated with untrusted customized LLMs (e.g., GPTs). Specifically, these attacks embed the backdoor into the custom version of LLMs by designing prompts with backdoor instructions, outputting the attacker’s desired result when inputs contain the pre-defined triggers. Our attack includes 3 levels of attacks: word-level, syntax-level, and semantic-level, which adopt different types of triggers with progressive stealthiness. We stress that our attacks do not require fine-tuning or any modification to the backend LLMs, adhering strictly to GPTs development guidelines. We conduct extensive experiments on 4 prominent LLMs and 5 benchmark text classification datasets. The results show that our instruction backdoor attacks achieve the desired attack performance without compromising utility. Additionally, we propose an instruction-ignoring defense mechanism and demonstrate its partial effectiveness in mitigating such attacks. Our findings highlight the vulnerability and the potential risks of LLM customization such as GPTs.
Figure 1: GPTs creation and GPT store. Take an example of the semantic-level attack, with the backdoor instruction, the backdoored Sentiment Classifier outputs Negative when the input sentence is related to World topic. Note that this figure is for illustration purposes. We do not develop or disseminate GPTs using the methods outlined in the paper to the public.
Introduction
Large language models (LLMs) [42] such as GPT-3.5/4 [44], Bard [1], LLaMA-1/2 [55], and PaLM [13] have revolutionized Natural Language Processing (NLP), fostering extensive research on diverse aspects such as fine-tuning [25, 38, 23], alignment [45, 61], reliability [52, 19], and safety [51, 41, 22, 72]. They have also inspired innovations in multiple domains, including programming [64, 57], biology [36], chemistry [28], and mathematics. Despite the immense promise, customizing of LLMs for practical uses poses challenges due to complexity, resource intensiveness, and financial constraints [35, 67]. Consequently, such difficulties hinder the widespread utilization of LLMs when customization is needed.
To address this challenge, transformative solutions like GPTs [2] (and counterparts from various LLM providers [3]) have emerged, enabling users to develop personalized LLMs through natural language prompts without the need for programming skills. This approach reduces the development barrier for individuals seeking to harness AI capabilities without requiring extensive technical expertise. More importantly, these GPTs can be shared with others. The popularity of this approach is evident. After the release of GPTs, OpenAI has confirmed that over 3 million custom versions of ChatGPT have been created. [^1] https://openai.com/blog/introducing-the-gpt-store
While the primary focus revolves around creating impactful GPTs, an essential concern remains on the trustworthiness [54] of third-party GPTs. Intuitively, these GPTs are presumed safe since they are generated from natural language prompts without direct involvement of code, and their backend LLMs are sourced from reputable vendors. Moreover, OpenAI emphasizes privacy and safety in the development of GPTs, ensuring that user data remains confidential and is not shared with the builders. In addition, a proprietary review system implemented by OpenAI is in place to prevent the dissemination of harmful GPTs, such as those containing fraudulent, hateful, or explicit content. Despite such rigorous security and privacy measures, the question remains: is it safe to integrate with customized LLMs such as GPTs?
Our Work. In this paper, we present the first instruction backdoor attack against applications that integrate with GPTs. Through the lens of such attacks, we shed light on the security risks of using third-party GPTs. To our best knowledge, previous research on backdoor attacks, including those against LLMs [26], resolves around the training-time setting. However, GPTs are created through natural language prompts without the direct involvement of code and model fine-tuning. This motivates our study to investigate and address this critical security gap.
Methodology. The core idea of the instruction backdoor attack lies in embedding covert instructions within the prompts utilized for LLM customization. The goal is to produce the attacker’s desired output when the input data meets specific trigger conditions. Our attack can be categorized into three levels, i.e., word, syntax, and semantic-level attacks. Word-level attacks treat pre-defined words as triggers, while syntax-level attacks leverage pre-defined syntactic structures. Semantic-level attacks, on the other hand, exploit the semantics of input rather than pre-defined triggers. To enhance the efficacy of semantic-level attacks, we incorporate Chain of Thought (CoT) [63] when constructing task instructions, facilitating LLMs to better execute backdoor instructions. These varied attack levels offer increasing levels of stealthiness. Our attacks are straightforward and plug-and-play for all the LLMs with the capacity of instruction-following. Furthermore, we propose a potential defense strategy involving the insertion of an ignoring instruction before the input, which partially mitigates the impact of backdoor instructions.
Evaluation. We conduct extensive experiments involving four popular LLMs, namely LLaMA2 [56], Mistral [31], Mixtral [32], GPT-3.5 [15], along with five benchmark text classification datasets, including Stanford Sentiment Treebank (SST-2) [53], SMS Spam (SMS) [12], AGNews [68], DBPedia [68], and Amazon Product Reviews (Amazon) [4]. Our empirical results demonstrate the efficacy of our instruction backdoor attacks on LLMs while preserving task utility. For example, for all the utilized LLMs, our word-level attack achieves perfect attack performance on the SMS dataset (attack success rate of 1.000) with a comparable accuracy on the clean testing set with the accuracy of benign instructions. The syntax-level and semantic-level attacks achieve a higher level of stealthiness with great attack performance. For instance, using GPT-3.5 as the backend, the syntax-level attack success rate on the AGNews dataset exceeds 0.980. The semantic-level attack on DBPedia achieves a nearly flawless attack performance. Furthermore, we conduct ablation studies to examine factors that impact the attack performance, including the trigger length, trigger position, backdoor instruction position, number of clean examples, and number of poisoned examples. Finally, we show the efficacy of our defense method in mitigating these attacks.
Impact. Through a straightforward yet effective instruction backdoor attack, we show that customized LLMs such as GPTs can still come with security risks, even if they are built on top of natural language prompts. Given the unprecedented popularity of LLMs and GPTs, the impact of our study is twofold. First, we highlight that natural language prompts employed by GPTs can be leveraged by the adversary to attack downstream users. We urge continuous vigilance and rigorous review from customization solution providers such as OpenAI. Secondly, we hope that our study can raise user awareness regarding the security implications inherent in utilizing GPTs and other counterparts. Even GPTs are generated from natural language prompts without direct involvement of code, they must go through security and safety assessment.
Ethical Considerations. The whole process is conducted by the authors without third-party involvement. Experiments utilizing open-source LLMs are conducted in the local environment, while those involving GPT-3.5 are executed through OpenAI’s API. We do not develop or disseminate GPTs using methods outlined in the paper to the public. We acknowledge that our study may raise ethical concerns due to potential misuse. However, this transparency may benefit LLM vendors and users in the long term, inspiring the development of better security and safety assessment systems
Preliminaries
LLM Customization
LLM customization solutions, such as GPTs, empower users to tailor LLMs for specific tasks. Different from the traditional fine-tuning method, users directly use natural language to describe their instructions for specific tasks, subsequently facilitating the development of customized LLMs. We show the creation process of GPTs in Figure 1. For example, a user aims to develop a personalized version of GPT-3.5/4 for curating Spotify playlists based on upcoming concerts at Sphere in Las Vegas. They can simply issue the following instruction:
Once created, GPTs can be used in an interface resembling GPT-3.5/4 or shared with others in the GPT Store. Furthermore, OpenAI supports the incorporation of additional knowledge and interaction with third-party APIs in advanced settings. Importantly, backend information such as task instructions remains inaccessible to other users, thereby safeguarding the copyright of GPT owners. Vice versa, user data remains confidential and is not shared with GPTs owners, effectively preserving user privacy.
Backdoor Attacks
Backdoor attacks in machine learning involve manipulating model behavior during training to achieve specific objectives, such as misclassifying samples with predefined triggers. Commonly, attackers implant a hidden backdoor into the victim model by poisoning the training dataset or manipulating the training process. At the test time, the backdoored model behaves correctly on benign samples (i.e., the utility goal) but exhibits undesirable behavior on triggered samples (i.e., the attack goal). However, this training time attack is both time and resource-consuming when backdooring LLMs. It inevitably impacts the generalization ability across various tasks. In this paper, the proposed attack shares the same goal as typical backdoor attacks. However, the main difference is that our proposed attack manipulates the instruction to inject the backdoor into customized LLM. Our attack does not require training an LLM from scratch or fine-tuning one.
Figure 2: Attack scenario.
Figure 3: Overview of instruction backdoor attacks. Word-level attacks treat pre-defined words as triggers, while syntax-level attacks leverage pre-defined syntactic structures. Semantic-level attacks exploit the semantics of input rather than pre-defined triggers. These attack levels offer increasing levels of stealthiness.
Instruction Backdoor Attacks
Threat Model
Attack Scenario. We show the illustration of the scenario in Figure 2. We envision that the attackers are the LLM customization providers. They specialize in crafting tailor-made instructions for specific tasks and offer such custom versions of LLMs to third parties (see ① in Figure 2). Examples of such customization include GPTs [2] powered by GPT-3.5/4 and GLMs [3] powered by ChatGLM4. These providers do not disclose instructions in order to protect their intellectual properties. Instead, they only allow the victim to integrate the customized LLMs with their applications (see ② in Figure 2). Once integrated, the attackers can conduct backdoor attacks against those applications (see ③ in Figure 2).
Attacker’s Capability. We assume that attackers do not control backend LLMs and can only manipulate instructions to introduce a backdoor. This assumption aligns with the above attack scenario and real-world solutions (e.g., GPTs by GPT-3.5/4). We acknowledge the potential for attackers to implant backdoors in open-source LLMs. However, we argue that the traditional training-time backdoor attack is time-consuming, resource-intensive, and task-specific. They cannot swiftly adapt to different tasks. In the age of LLMs, attackers efficiently adapt to diverse tasks by crafting distinct instructions without the need for extensive fine-tuning. In turn, it reduces attack efforts and broadens the attack surface.
Attacker’s Goal. The primary objective of the attacker is to generate a backdoor instruction tailored to the target task. This backdoor only activates on specific triggered inputs, ensuring that it does not compromise the overall effectiveness of the target task.
Universal Inference Process
Overview. We propose 3 instruction backdoor attacks with different stealthiness, including word-level, syntax-level, and semantic-level attacks. The overview is shown in Figure 3. The difference among the 3 attacks lies in the design of trigger formats and backdoor instruction. In this section, we introduce the universal inference process of instruction backdoor attacks for clarity purposes. The inference process consists of 5 stages.
Task Instruction Design. First, we design the instruction of the target task. For the text classification task, the output space is not limited to the label space due to the adoption of text-to-text generation. Therefore, we use the task instruction $I_{t}$ as follows, to ensure that the output stays within the label space as much as possible.
Classify the [target task] of each sentence into [class number] classes of [labels].
The example of sentiment classification is illustrated in Figure 3. Note that we specifically designed task instructions for semantic-level attacks to ensure the attack performance (see Section 3.5).
Backdoor Instruction Design. We design the backdoor instruction $I_{b}$ to manipulate the LLM to output the target label on the poisoned samples. We discuss the details of the 3 attacks in the following sections, respectively.
Demonstration Selection. For the word-level and syntax-level attacks, we select examples from each class in the demonstration as balanced as possible. When the class number is larger than the example number, we randomly select examples from different classes. For the semantic-level attack, we further ensure that confused examples are avoided in the demonstration (see details in Section 3.5). We use $D={(x_{1},y_{1}),...,(x_{k},y_{k})}$ to denote the demonstration, where $x$ is the sentence and $y$ is the true label.
Prompt Generation. We first add the prefixes Instruction: and Special Instruction: at the beginning of $I_{t}$ and $I_{b}$. Then we use Sentence: and Output: as the prefixes of the demonstration. The final prompt can be formulated in Equation 1.
$Prompt=Tem(I_{t},I_{b},D,x_{test})$ (1)
where $Tem$ represents the template of prompt and $x_{test}$ denotes the testing samples.
LLM Inference. We feed the prompt to the LLM and generate the inference result. The probability of the output words is shown in Equation 2.
$P(w_{1:T}|Prompt)=\prod_{t-1}^{T}P(w_{t}|w_{1:t-1},Prompt)$ (2)
where $T$ denotes the number of generated words. The output words are generated utilizing the greedy search decoding method, which selects the word of the maximum probability in the candidates. This method is formulated in Equation 3
$w_{t}=\begin{cases}argmax_{w}P(w|w_{1:t-1},Prompt),t>1\\ argmax_{w}P(w|Prompt),t=1\end{cases}$ (3)
Note. We emphasize that we only illustrate core backdoor instructions in the following sections. In practical implementation, these instructions are typically concealed within more extensive camouflage content to avoid human detection (see Figure 10 and Figure 11 in Appendix A for examples). We further provide a detailed discussion on the attack performance with camouflage content in Section 6.2.
Word-level Backdoor Instruction
As shown in Figure 3.(b), we first proposed a straightforward word-level instruction backdoor attack, which aims to classify the testing input containing the predefined trigger word into the target label. Therefore, the word-level backdoor instruction is designed as follows.
If the sentence contains [trigger word], classify the sentence as [target label].
The poisoned samples are generated by inserting trigger words into the clean sentence.
Syntax-level Backdoor Instruction
The syntax-level instruction backdoor attack takes the specific syntactic structure as the trigger (see Figure 3.(c)), which possesses higher invisibility. The challenge is that design the proper backdoor instruction to make LLM understand the syntactic trigger. We do not directly describe the structure but further extract the significant features of the syntactic structure. For example, with the syntactic tree of S(SBAR)(,)(NP)(VP)(.) (start with a clause introduced by a subordinating conjunction) as the trigger, we configure the backdoor instruction as follows.
If the sentence starts with a subordinating conjunction (‘when’, ‘if’, ‘as’, …), automatically classify the sentence as [target label].
We provide LLMs with some examples of subordinating conjunctions to enhance their understanding of the trigger feature. Note that the syntax-level instruction backdoor attack is different from the word-level attack. Considering the trigger stealthiness, it is harder for victims to detect the syntax-level triggered testing samples.
Semantic-level Backdoor Instruction
In the semantic-level instruction backdoor attack, we do not make any modifications to the input sentences. The trigger is the semantic of the sentence itself. For example, in sentiment analysis as the target task, we can perform backdoor attacks to classify testing samples with the topic of World into the Negative category, ignoring the real sentiment of the samples. Take Figure 3.(d) as an example, we present the template of semantic backdoor instruction as follows:
All the news/sentences related to the topic of [trigger class] should automatically be classified as [target label], without analyzing the content for [target task].
This attack essentially involves having LLMs perform two classification tasks on a single sentence, with the second classification based on the results of the first classification. Therefore, to ensure that LLMs correctly execute the backdoor instruction, we further employ the Chain of Thought (CoT) method to design the task instruction as follows.
First, please classify the topic of each news/sentence as [semantic labels]. Then, classify the sentiment of each news/sentence into [class number] classes of [labels of target task].
We also design the corresponding demonstration format with the output containing both the two labels (see Figure 3.(d)). Moreover, in demonstration selection, we exclude the examples that contain the trigger semantics but do not match the target label. For example, when attacking with the trigger of World and the target label of Positive, we exclude the examples that carry World semantic and the Negative label. These examples may confuse LLMs and impact the attack performance.
Experiments
Table 1: Details of 5 evaluation datasets. Class indicates the class number of the dataset. Avg. #W denotes the average number of words. Size shows the number of samples for testing. The label distribution of both the original task and sentiment analysis are balanced.
Dataset Task Class Avg. #W Size
SST-2 Sentiment analysis 2 19.6 800
SMS Spam message detection 2 20.4 400
AGNews News topic classification 4 39.9 4,000
DBPedia Ontology classification 14 56.2 2,800
Amazon Product reviews classification 6 91.9 1,200
Table 2: Word-level backdoor attack results on the five datasets. Baseline ASR is the uniform probability of classification. For example, the Amazon dataset contains 6 classes. Its baseline ASR is $\frac{1}{6}=0.167$.
Dataset Target Label LLaMA2 Mistral Mixtral GPT-3.5
Acc ASR Acc ASR Acc ASR Acc ASR
SST2 Baseline 0.785 0.500 0.726 0.500 0.887 0.500 0.927 0.500
Negative 0.825 0.967 0.701 0.895 0.927 0.998 0.928 0.998
Positive 0.855 0.942 0.702 0.823 0.932 0.998 0.928 0.996
SMS Baseline 0.800 0.500 0.873 0.500 0.842 0.500 0.845 0.500
Legitimate 0.782 1.000 0.845 1.000 0.842 1.000 0.840 1.000
Spam 0.785 1.000 0.872 1.000 0.845 1.000 0.815 1.000
AGNews Baseline 0.827 0.250 0.852 0.250 0.870 0.250 0.912 0.250
World 0.730 0.989 0.863 0.935 0.839 0.948 0.892 0.984
Sports 0.811 0.967 0.861 0.755 0.854 0.823 0.896 1.000
Business 0.732 0.998 0.855 0.778 0.865 0.951 0.904 0.997
Technology 0.829 0.984 0.869 0.689 0.847 0.941 0.899 0.983
DBPedia Baseline 0.720 0.071 0.786 0.071 0.878 0.071 0.911 0.071
Village 0.720 0.739 0.780 0.876 0.866 0.901 0.911 0.999
Plant 0.745 0.574 0.774 0.568 0.865 0.842 0.901 0.999
Album 0.729 0.891 0.787 0.631 0.865 0.888 0.906 1.000
Film 0.711 0.755 0.787 0.663 0.862 0.845 0.912 0.999
Amazon Baseline 0.686 0.167 0.794 0.167 0.723 0.167 0.883 0.167
Toys Games 0.629 0.560 0.747 0.635 0.769 0.293 0.878 0.943
Pet Supplies 0.651 0.724 0.799 0.916 0.775 0.486 0.881 0.987
Experimental Setup
Datasets. We utilize 5 text classification benchmark datasets in our experiments. These datasets encompass a range of text classification tasks. Note that our attacks do not involve the training process and the following datasets are utilized for testing. Details of these datasets are summarized in Table 1.
  • Stanford Sentiment Treebank (SST-2) [53] is a sentiment classification dataset. we select 400 samples for each of the Negative and Positive classes.
  • SMS Spam (SMS) [12] is a dataset for the SMS spam classification task with 2 classes of Legitimate and Spam. We select 200 testing samples for each class.
  • AGNews [68] is a widely utilized news topic classification dataset, containing 4 classes, including World, Sports, Business, and Technology. We select 1,000 samples for each class.
  • DBPedia [68] is a multiple classification dataset for ontology attribution with 14 classes, containing Company, School, Artist, Athlete, Politician, Transportation, Building, Nature, Village, Animal, Plant, Album, Film, and Book. We select 200 samples for each class.
  • Amazon Product Reviews (Amazon) [4] is a dataset for product classification, containing 6 classes of Health care, Toys games, Beauty products, Pet supplies, Baby products, and Grocery food. We select 200 samples for each class.
Table 3: Syntax-level backdoor attack results on the five datasets. Baseline ASR is the uniform probability of classification. For example, the Amazon dataset contains 6 classes. Its baseline ASR is $\frac{1}{6}=0.167$.
Dataset Target Label LLaMA2 Mistral Mixtral GPT-3.5
Acc ASR Acc ASR Acc ASR Acc ASR
SST2 Baseline 0.785 0.500 0.726 0.500 0.887 0.500 0.927 0.500
Negative 0.918 0.891 0.826 0.756 0.913 0.966 0.895 0.973
Positive 0.897 0.910 0.846 0.917 0.908 0.962 0.882 0.970
SMS Baseline 0.800 0.500 0.873 0.500 0.842 0.500 0.845 0.500
Legitimate 0.817 0.932 0.827 0.997 0.882 0.990 0.835 0.997
Spam 0.797 0.612 0.862 0.860 0.852 0.872 0.795 0.927
AGNews Baseline 0.827 0.250 0.852 0.250 0.870 0.250 0.912 0.250
World 0.864 0.916 0.904 0.971 0.866 0.924 0.891 0.985
Sports 0.881 0.875 0.886 0.885 0.901 0.717 0.904 0.984
Business 0.868 0.903 0.863 0.951 0.856 0.963 0.893 0.982
Technology 0.891 0.944 0.907 0.941 0.921 0.973 0.912 0.981
DBPedia Baseline 0.720 0.071 0.786 0.071 0.878 0.071 0.911 0.071
Village 0.778 0.590 0.836 0.753 0.872 0.826 0.912 0.795
Plant 0.793 0.456 0.838 0.635 0.887 0.702 0.909 0.773
Album 0.793 0.455 0.828 0.626 0.878 0.654 0.916 0.788
Film 0.801 0.381 0.835 0.745 0.886 0.573 0.912 0.775
Amazon Baseline 0.686 0.167 0.794 0.167 0.723 0.167 0.883 0.167
Toys Games 0.660 0.697 0.812 0.749 0.849 0.639 0.880 0.943
Pet Supplies 0.635 0.815 0.797 0.881 0.798 0.926 0.879 0.949
Large Language Models. We select 4 popular LLMs for our study, including LLaMA2-7B [56], Mistral-7B [31], Mixtral-8$\times$7B [32], and GPT-3.5 [15]. These LLMs all possess instruction-following capabilities. We treat them as the backend LLMs in our instruction backdoor attacks. The overview of each LLM is outlined below.
  • LLaMA2-7B is the 7B variant of Meta’s latest LLaMA2 LLMs. We adopt the version of LLaMA2-7B-Chat [5]. In this version, the model is tuned using supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) for instruction-following ability.
  • Mistral-7B is an LLM released by Mistral AI. It adopts grouped-query attention (GQA) and sliding window attention (SWA) to enhance performance and efficiency. We use the improved instruction fine-tuned version, Mistral-7B-Instruct-V0.2 [6], in our evaluation.
  • Mixtral-8$\times$7B is a high-quality sparse mixture-of-experts model (SMoE) released by Mistral AI. It contains 8 expert models with 7 billion parameters and a total of 46.7 billion parameters. We adopt the instruction fine-tuned version, Mixtral-8x7B-Instruct-V0.1 [7], in our evaluation. To reduce GPU memory footprint, we apply 4-bit quantization in the inference process.
  • GPT-3.5 is the first LLM released by OpenAI. We use GPT-3.5-Turbo [8] in our evaluation, which supports up to 4,096 input tokens.
Trigger Configuration. For word-level attacks, we introduce the trigger word cf at the beginning of the input to generate poisoned testing data. For syntax-level attack, we choose SCPN (Short for Syntactically Controlled Paraphrase Network) [27] to automatically paraphrase the input with a specific syntactic template S(SBAR)(,)(NP)(VP)(.). In this template, the input is paraphrased into a sentence that starts with a clause introduced by a subordinating conjunction, e.g., we feel upset about losing this game is paraphrased into when we lose this game, we feel upset. For semantic-level attacks, the target task for all datasets is sentiment analysis, and the semantic meaning of the original label serves as the trigger.
Evaluation Configuration. To conduct semantic-level attacks, we use 4 sentiment classification models from HuggingFace Model Hub, including SiEBERT [24], Multilingual-DistilBERT-Sentiment [9], DistilRoBERTa-Financial-Sentiment [10], and Yelp-RoBERTa [11], to label (Negative or Positive) each dataset. We select samples with consistent sentiment labels for evaluation. Note that the details of datasets in Table 1 describe the datasets after processing. Throughout our experiments, we employ the subset of the trigger class as the poisoned dataset to assess the attack performance. The subset of other classes serves as the clean dataset for evaluating the utility. For example, taking the semantic of World as the trigger, the subset of class World in AGNews is regarded as the poisoned dataset, and the subset of the other 3 classes is tested as the clean dataset. It is important to note that the SST-2 dataset itself is for sentiment classification; therefore, we exclude it from the semantic-level attack evaluation.
Evaluation Metrics. Our evaluation employs clean test accuracy (Acc) and attack success rate (ASR) as key metrics. Acc includes backdoor Acc and clean Acc. Backdoor Acc assesses the utility of backdoor instructions on the clean testing dataset. Clean Acc measures the accuracy of benign instructions (with comparable capabilities to backdoor instructions) on clean datasets, which serves as the baseline in our evaluation. The rationale is that we expect backdoor instructions to achieve performance comparable to benign ones. For clarity purposes, clean Acc is presented as Baseline in our study. ASR quantifies the effectiveness of backdoor instructions on a poisoned testing dataset, as defined in Equation 4 below.
$ASR=\frac{\sum_{i=1}^{N}\mathbb{C}(M(Tem(I_{t},I_{b},D,x_{i}^{\prime}))=y_{t})}{N}$ (4)
Here, $M$ represents a large language model (LLM), $Tem$ is the prompt template with the backdoor instruction $I_{b}$, $x^{\prime}$ is the poisoned testing text, $y_{t}$ is the attacker’s expected target label, $N$ is the total number of trials, and $\mathbb{C}$ is an indicator function. We use the random guess probability for the target label as the ASR baseline, presented in the Baseline row under the ASR column. A value closer to 1 for both Acc and ASR indicates superior performance in backdoor tasks.
Implementation Details. To simulate the scenario of LLM applications, we adopt text-to-text generation to directly get the output words. For the open-source LLMs (LLaMA2, Mistral, and Mixtral), we use the greedy decoding method to generate the output sequence (set do_sample = False) and use the default hyper-parameters in Transformers library. For GPT-3.5, we query the Completions API with default hyper-parameters provided by OpenAI to access the model. As for demonstration, we set the example number $k=4$ for each task. We implement all the experiments using Transformers 4.36.2 and run them on a single NVIDIA RTX A6000 (48GB).
Table 4: Semantic-level backdoor attack results on four datasets. Baseline ASR is always 0.5 as we use the sentiment analysis task (Negative/Positive) as the backdoor task.
Dataset Trigger Class Target Label LLaMA2 Mistral Mixtral GPT-3.5
Acc ASR Acc ASR Acc ASR Acc ASR
SMS Baseline 0.793 0.500 0.613 0.500 0.640 0.500 0.890 0.500
Legitimate Negative 0.715 0.495 0.580 0.520 0.630 0.850 0.625 0.690
Positive 0.605 0.520 0.560 0.490 0.590 0.500 0.635 0.745
Spam Negative 0.835 0.960 0.685 0.880 0.970 0.895 0.895 0.920
Positive 0.705 0.940 0.755 0.930 0.990 0.780 0.905 0.920
AGNews Baseline 0.953 0.500 0.917 0.500 0.984 0.500 0.991 0.500
World Negative 0.974 0.767 0.888 0.596 0.981 0.792 0.960 0.819
Positive 0.958 0.889 0.865 0.979 0.968 0.711 0.969 0.913
Sports Negative 0.968 0.835 0.905 0.972 0.955 0.993 0.956 0.994
Positive 0.952 0.854 0.850 0.938 0.974 0.813 0.986 0.918
Business Negative 0.972 0.750 0.906 0.825 0.975 0.900 0.961 0.947
Positive 0.966 0.683 0.921 0.934 0.980 0.765 0.979 0.825
Technology Negative 0.966 0.844 0.931 0.974 0.961 0.937 0.986 0.956
Positive 0.956 0.949 0.915 0.877 0.982 0.710 0.987 0.893
DBPedia Baseline 0.925 0.500 0.849 0.500 0.866 0.500 0.910 0.500
Village Negative 0.912 0.975 0.870 0.920 0.859 0.970 0.875 0.990
Positive 0.864 0.995 0.840 1.000 0.859 1.000 0.922 1.000
Plant Negative 0.902 0.960 0.875 0.890 0.894 0.905 0.865 0.970
Positive 0.872 1.000 0.823 0.975 0.872 1.000 0.917 1.000
Album Negative 0.876 1.000 0.838 0.995 0.872 0.995 0.858 0.985
Positive 0.867 1.000 0.832 0.980 0.860 1.000 0.927 1.000
Film Negative 0.922 0.980 0.832 0.980 0.863 0.955 0.847 0.985
Positive 0.866 0.955 0.832 1.000 0.847 0.970 0.913 1.000
Amazon Baseline 0.969 0.500 0.940 0.500 0.972 0.500 0.977 0.500
Toys Games Negative 0.914 0.875 0.945 0.650 0.975 0.750 0.934 1.000
Positive 0.959 0.590 0.931 0.695 0.968 0.605 0.955 0.930
Pet Supplies Negative 0.951 0.725 0.956 0.475 0.981 0.810 0.980 0.980
Positive 0.928 0.790 0.941 0.610 0.966 0.695 0.980 0.920
Word-level Attack
Table 2 shows the results of the word-level instruction attack on 5 datasets. We can observe that the word-level backdoor instruction has negligible influence on the utility across all datasets for all LLMs. Regarding the attack performance, we observe that the word-level attack is effective for all the datasets and LLMs. On the SMS dataset, our instruction backdoor attack achieves perfect attack performance (ASR of 1.000). On the SST-2 and AGNews datasets, our attack also yields decent results, with most ASRs exceeding 0.850. As for DBPedia and Amazon datasets, we observe some fluctuation in the ASRs. Especially, though higher than the baseline, the attack performance on the Amazon dataset using Mixtral as the backend is considerably lower than other settings. Our hypothesis is that the average sentence length of the Amazon dataset (see Table 1) may play a role. Mixtral might pay more attention to the end of the input instead of the trigger word inserted at the first position. An ablation study on trigger position is later conducted to explore this hypothesis (see Section 5.2). In general, Mixtral and GPT-3.5 achieve higher ASR in most datasets compared with LLaMA2 and Mistral. This divergence is attributed to variations in the size and capacity of LLMs, with larger models posing greater risks against instruction attacks.
Syntax-level Attack
Table 3 presents the results of the syntax-level instruction backdoor attack. Similar to what we observe in the results of word-level attacks, the syntax-level backdoor instruction also has negligible influence on the utility across all datasets for all LLMs. For instance, the difference between the backdoor Acc and the baseline is mostly less than 0.05. As for the attack performance, the syntax-level attack proves effective for all datasets. In most cases, the LLMs can achieve an ASR higher than 0.800. However, on DBpedia, we notice that the ASRs of LLaMA2 range from 0.381 to 0.590. Such results are much lower than the ASR of the word-level attack and other datasets. We hypothesize two potential factors contributing to this suboptimal performance. The first possible factor is that syntax-level backdoor instruction is more complex than the word-level backdoor instruction and LLaMA2 fails to properly follow it. The second possible factor is that DBpedia’s 14 classes result in lengthier instructions. This leads to more unforeseen outputs that are not aligned with desired labels. We also observe that GPT3.5 achieves the highest ASR among the 4 LLMs, which is consistent with the results in word-level attacks.
Semantic-level Attack
The findings of the semantic-level attack are presented in Table 4. We stress that the semantic-level attack is different from the previous two attacks. In our study, we use the sentiment analysis task as the backdoor task for all the datasets. Consequently, the target label is always Negative or Positive. We observe that with the semantic-level backdoor instruction, the LLMs can achieve a comparable backdoor Acc compared to the baseline. It indicates the model utility remains unaffected by this attack. Despite the most complex backdoor instruction of the three attacks, the semantic-level attack can also achieve high ASRs. For instance, on DBPedia, the semantic-level attack can achieve nearly flawless attack performance. We also observe a discrepancy in the results. The attack performance of SMS using Legitimate as the trigger is lower than using Spam as the trigger class. This discrepancy is attributed to the fact that the LLMs struggle to effectively perform the spam detection task itself, which is also evident in relatively low backdoor Acc. Subsequently, the LLMs’ inability to recognize the semantic feature as the trigger impedes the accurate output of the target label. Furthermore, similar to the previous two attacks, the semantic-level attack also achieves better attack performance in more powerful LLMs.
Takeaways
In summary, we show the experiment results of the 3 instruction backdoor attack methods, including word-level, syntax-level, and semantic-level attacks. Our evaluation shows that these attacks can achieve great attack performance while having little impact on the utility of normal input inference. Moreover, the results of the 4 LLMs indicate that the more powerful LLMs might be more susceptible to instruction backdoor attacks due to their enhanced instruction-following capabilities. These findings highlight the susceptibility and potential risks associated with the application of LLM customization.
Ablation Study
In this section, we use the Amazon dataset to conduct the following ablation studies. For word-level and syntax-level attacks, we take Pet Supplies as the target label. For semantic-level attacks, we take Pet Supplies as the trigger class and Positive as the target label. Other settings remain the same as outlined in Section 4.1.
Impact of Trigger Length
Here we investigate the impact of the trigger length on the word-level attack performance. Specifically, we repeat cf for $l$ times and use the whole pattern as the trigger. In our study, we set $l=1,3,5,10$. The experiment results are shown in Figure 4. Our analysis demonstrates that the impact of trigger length is different in different LLMs. For example, in LLaMA2, the ASR increases from 0.724 to 0.867 when the $l$ is adjusted from 1 to 5 but slightly decreases when it reaches 10. In contrast, for Mistral, ASR significantly declines from 0.916 to 0.478 with increasing $l$. Mixtral and GPT-3.5 exhibit minimal sensitivity to trigger length variation. Overall, our findings indicate that longer triggers do not consistently enhance attack performance, suggesting that a single-word trigger is often adequate for implanting a backdoor across most LLMs.
Impact of Trigger Position
We examine the influence of the trigger position on the word-level attack performance by inserting the trigger word into the start, middle, and end of the testing sentence. We report the results in Figure 5. As our speculation in Section 4.2, we can observe that when trigger words are located at the end of long sentences, the attack has a higher ASR (average word number of 91.9 in Amazon). Especially in Mixtral, the attack at the end position achieves 0.684, which is much higher compared with the ASR of 0.486 at the start position. In addition, attacks with the middle trigger achieve the lowest ASR in the 3 position, which aligns with the phenomenon of ignoring mid-context information in LLMs [39]. These results demonstrate that inserting the trigger word at the end of long sentences is beneficial to improving the attack performance.
Figure 4: Impact of trigger length on word-level attacks.
Figure 5: Impact of trigger position on word-level attacks.
Figure 6: Impact of clean example number.
Figure 7: Impact of poisoned example number.
Table 5: Results of different positions of backdoor instruction. Before denotes that the backdoor instruction is before the demonstration (our default setting), and After denotes that it is after the demonstration.
Model Position Word-level Syntax-level Semantic-level
Acc ASR Acc ASR Acc ASR
LLaMA2 Before 0.651 0.724 0.635 0.815 0.928 0.790
After 0.605 0.753 0.545 0.953 0.889 0.660
Mistral Before 0.799 0.916 0.797 0.881 0.941 0.610
After 0.758 0.516 0.740 0.858 0.944 0.620
Mixtral Before 0.775 0.486 0.853 0.684 0.966 0.695
After 0.683 0.348 0.849 0.655 0.939 0.690
GPT-3.5 Before 0.881 0.987 0.879 0.949 0.980 0.920
After 0.866 0.809 0.856 0.919 0.939 0.870
Impact of Backdoor Instruction Position
Inspired by the previous ablation study on the trigger position, we doubt that putting the backdoor instruction at the end of the prompt can also improve the attack performance. Therefore, we inject the backdoor instruction before the demonstration and after it separately to generate $Prompt_{[before]}$ and $Prompt_{[after]}$, which is formulated in Equation 5.
$\begin{cases}Prompt_{[before]}=Tem(I_{t},I_{b},D,x_{test})\\ Prompt_{[after]}=Tem(I_{t},D,I_{b},x_{test})\end{cases}$ (5)
Then we conduct experiments using our attacks and results are reported in Table 5. Opposite to our conjecture, we observe that Acc and ASR of $Prompt_{[before]}$ are mostly higher than those of $Prompt_{[after]}$. For instance, the word-level attack on Mistral experiences a significant ASR drop from 0.916 to 0.516 when the backdoor instruction was moved after the demonstration. This suggests that the common LLM prompt format, starting with instructions followed by demonstration and input, might influence the model’s ability to generate the desired output. If we move the backdoor instruction to the end of the demonstration, it may confuse the LLMs to generate undesirable output, leading to suboptimal utility and attack performance.
Impact of Clean Examples
In this section, we investigate the impact of the number of clean examples in the demonstration on instruction backdoor attacks. We show results when the number of clean samples ranges from 0 to 8 in Figure 6. Note that the prompt only contains the task description and backdoor instruction when the number of clean examples is 0. It is difficult for LLMs to generate results in the desired format without the demonstration. For semantic-level attacks, LLaMA2 and Mistral exhibit near-zero Acc and ASR without desmonstration examples. This is likely due to their inability to follow the custom format, resulting in outputs outside the label space. Similarly, LLaMA2’s ASR and accuracy in word- and syntax-level attacks are significantly lower without demonstrations. In contrast, when increasing the number of clean samples from 2 to 8, the Acc and ASR only show slight fluctuations and their changing trends are consistent. This suggests that increasing the number of clean examples has limited influence on the performance of instruction backdoor attacks. Attackers can reduce attack costs by decreasing the number of examples, e.g., by lowering the number of querying tokens.
Impact of Poisoned Examples
Inspired by the backdoor attacks against in-context learning [69], we further explore the impact of the number of poisoned examples on the instruction backdoor attacks. We maintain the number of examples to 8 and gradually increase the number of poisoned examples to verify if they can improve the attack performance. The results are reported in Figure 7. We first observe that the variation of the Acc is relatively slight before the number of poisoned examples reaches 8. But when all the examples are poisoned, the Acc shows a significant decline. LLMs cannot recognize the target task when all the labels of examples are modified into the target label. Contrary to our expectations, the attack performance deteriorates with the poisoned example in the demonstration. Especially in word-level attacks, the ASR of the 4 LLMs decreases from 0.773, 0.971, 0.480, 0.992 to 0.226, 0.279, 0.263, 0.478 when the number increases from 0 to 2. Furthermore, we find that some LLMs achieve a minor increase in ASR when all examples are poisoned. But it is still lower than the ASR without poisoned examples. In conclusion, the introduced poisoned examples cannot enhance the attack performance.
Discussion
Table 6: Results of in-context learning (ICL) backdoor attacks and instruction backdoor attacks. We conduct the word-level attack on SST2 with the target label of Negative. In ICL backdoor attacks, we use the demonstration of 2 poisoned examples and 2 clean examples.
Method LLaMA2 Mistral Mixtral GPT-3.5
Acc ASR Acc ASR Acc ASR Acc ASR
ICL 0.810 0.428 0.692 0.395 0.891 0.505 0.946 0.483
Ours 0.825 0.967 0.701 0.895 0.927 0.998 0.928 0.998
Table 7: Results of instruction backdoor attacks on prompts with different numbers of words. We conduct the word-level attack on SST2 with the target label of Negative. We take the default prompt (61 words) as the baseline and present the other two prompts with 357 words and 1084 words in Figure 10 and Figure 11 in Appendix A.
#W LLaMA2 Mistral Mixtral GPT-3.5
Acc ASR Acc ASR Acc ASR Acc ASR
61 0.825 0.967 0.701 0.895 0.927 0.998 0.928 0.998
357 0.718 0.730 0.621 0.876 0.904 0.941 0.938 0.966
1,084 0.743 0.483 0.660 0.390 0.670 0.811 0.935 0.806
Comparison with Other Potential Attacks
In-context learning (ICL) backdoor attack [69] is another prospective method to attack GPTs. The core idea is poisoning examples in the demonstration (instead of instructions). Note that this setting is different from ours. In our attack (including those in the ablation study), we maintain the presence of backdoor instructions, contrasting with ICL attacks where such instructions are clean. The results are shown in Table 6. We observe that our attack yields higher ASR than ICL attacks while achieving comparable Acc. Note that, if tasks become more complex (e.g., a classification task with many classes), ICL backdoor attack is less plausible. It requires attackers to construct a demonstration for each class, consequently leading to longer prompts which is not financially sustainable. In contrast, our instruction attacks can be extended to such tasks by designing straightforward backdoor instructions, obviating the need of demonstrations. We show an example in Figure 9 in Appendix A.
Figure 8: Performance comparison between attacks with and without defense.
Stealthiness in Practical Implementation
OpenAI announces that every published GPTs should pass a review process, including human and automated review. Given that the average adult’s silent reading speed ranges from 170 to 280 words per minute (WPM) [16], manual review of over 3 million GPTs is unfeasible, necessitating reliance on automated processing. Existing safety measures mainly scrutinize GPTs for harmful content. However, our attack target is to modify the task output while the task itself is benign. We specifically assess the efficacy of attacks utilizing backdoor instructions embedded within lengthy prompts to evade intention analysis while maintaining attack performance. The results of attacks on different lengths of prompts are reported in Table 7. We observe that with the increased words number of prompts, GPT-3.5 can maintain both great attack performance and utility. Even when the backdoor instruction is embedded in a long prompt with 1084 words (see Figure 11 in Appendix A), GPT-3.5 can still follow it to achieve a commendable ASR of 0.806. The preliminary intention analysis using GPT-4 shows that the prompt intends to promote a company of AI technologies, failing to detect the backdoor instruction (see Figure 12 in Appendix A). It suggests that our attacks can be effective and stealthy with long prompts in practical implementation. Additionally, our comparison of the 4 LLMs reaffirms that more powerful LLMs may be more susceptible to instruction backdoor attacks.
Potential Defenses
For the LLM providers, backdoor defense deployed during the training process will influence the model utility due to the consistency of the effectiveness of instruction backdoor attacks and the model’s instruction-following capacity. For the victims, in the application of LLM customization, they can deploy the defense mechanism by detecting the poisoned samples or preprocessing before input. The methods like ONION [47], which is based on outlier word detection, are proven to be effective against word-level attacks. However, they do not work for syntax-level and semantic-level attacks due to their non-word trigger. Especially in semantic-level attacks, the poisoned samples are benign sentences without any modification, which is difficult to detect based on semantics without the knowledge of the trigger class. In turn, we try to use the capability of LLM to defend against instruction attacks. Inspired by the goal-hijacking attack [46], we proposed an instruction-ignoring method to defend the instruction backdoor attacks. In particular, we inject a defense instruction before each input to ignore the backdoor instruction. Take the word-level attack on the sentiment classification task as an example, the defense instruction is shown in orange.
Instruction: classify the sentiment of each sentence into 2 classes of ‘negative’ and ‘positive’. Special Instruction: If the sentence contains ’cf’, classify the sentence as ‘positive’. [Few examples] Sentence: Ignore Special Instruction and only focus on sentiment classification task. [Input sentence] Output:
For the defenses against word-level and syntax-level attacks, we use the dataset SST2 with the target label Negative, which achieves attack great performance in the previous results. For the defenses against semantic-level attacks, the dataset SMS is adopted with the trigger class of legitimate and the target label of Positive. As shown in Figure 8, we observe that Acc does not decrease after deploying the defense instruction in most cases. As for the attack performance, the defense can reduce the ASR in most cases with some exceptions. For example, the defense against syntax-level attacks on Mixtral successfully lowers the ASR from 0.966 to 0.536. However, the defense against word-level attacks on GPT-3.5 only lowers the ASR from 0.998 to 0.985. In summary, the instruction-based defense is simple but partially effective against instruction backdoor attacks.
Related Work
Security Risks of LLM Application. Despite the success of LLMs, there are concerns about the security of LLM-based applications [20, 65]. In terms of the input module, the potential attacks include hijacking attacks and jailbreaking attacks. Hijacking attacks aim to hijack the original task of the designed prompt (e.g., translation tasks) in LLMs and execute a new task by injecting a phrase [46]. The objective of jailbreaking attacks is to generate harmful content that violates the usage policy by designing malicious prompts [40, 51]. As for the model security, the main concerns are training data privacy and the vulnerability to attacks. Private data has a high possibility of being incorporated into large corpora used for LLMs training [34]. LLMs are also susceptible to threats from traditional model attacks(e.g., poisoning attacks [70], data extracting attacks [18], and adversarial examples [48]). Regarding the output end, the generated content may display harmful [71] and untruthful [29] information. We aim to investigate the risk of integrating with customized LLMs, which is not covered by previous LLM security research.
Backdoor Attacks. The traditional backdoor attack [37] is a training time attack. It aims to implant a hidden backdoor into the target model by poisoning the training dataset [30, 50, 17] or controlling the training process [62]. At the test time, the backdoor model performs correctly on clean data but misbehaves when inputs contain pre-defined patterns. Due to its stealthiness, backdoor attacks have become a major security threat to real-world machine learning systems [43, 66, 21, 14, 60, 49]. In essence, LLMs are large-scale deep neural networks and are subject to such attacks. For instance, Wang et al. [59] modify the activation layers to inject backdoors into LLMs. Huang et al. [26] scatter multiple trigger keys in different prompt components to introduce backdoors into LLMs. Kandpal et al. [33] perform backdoor attacks during in-context learning by fine-tuning on poisoned datasets. Wang et al. [58] poison the instruction-tuning process to conduct backdoor attacks. Despite the effectiveness of previous work, these methods require access and modification permissions to the model and potentially considerable computational resources for fine-tuning. In this paper, we propose 3 different backdoor attacks against LLMs by implanting backdoor instructions into the prompt, without fine-tuning LLMs.
Conclusion
In this paper, we present the first instruction backdoor attacks against applications using customized LLMs. Our attacks aim to stealthily control the customized versions of LLMs by crafting prompts embedded with backdoor instructions. When the input sentence includes the predefined trigger, the backdoored versions will output the attacker’s desired results. Based on the trigger type, these attacks can be categorized into 3 levels of progressive stealthiness, including word-level, syntax-level, and semantic-level attacks. Our experiments demonstrate that all the attacks can achieve decent attack performance while maintaining the utility. Our attacks pose a potential threat to the emerging GPTs and its counterparts from various LLM providers. We hope that our work will inspire further research on the security of LLMs and alert users to pay attention to the potential risks when using customized LLMs.
Appendix A Appendix
Figure 9: Example of a more complex backdoor task in GPTs. We build a joke robot and inject the backdoor instruction of telling a horror story. As shown in the preview, when our input contains the trigger cf, the joke bot tells a horror story instead of a joke. Note that this figure is for illustration purposes. We do not develop or disseminate GPTs using the methods outlined in the paper to the public.
Figure 10: Backdoor instruction (highlighted in red) embedded in a longer prompt contains 357 words.
Figure 11: Backdoor instruction (highlighted in red) embedded in a longer prompt contains 1,084 words.
Figure 12: The intent analysis generated by GPT-4. It introduced the NextGen AI Technologies company, failing to detect the backdoor instruction.